OpenStudy (anonymous):

help me prove this plz use only sin^2A+cos^2A=1formet sec^4A-1=2tan^2A+tan^4A

OpenStudy (anonymous):

Well with that format we can't use this for sec and tan

OpenStudy (anonymous):

By that I mean, we should use this formula using your formula \[\Large \sin^2+\cos^2=1\] Divide all sides by cos and you get \[\Large {\sin^2 \over \cos^2} +\frac{\cos^2}{\cos^2}=\frac{1}{\cos^2}\] \[\Large \tan^2+1=\sec^2\]

OpenStudy (anonymous):

rearranging we get \[\Large \tan^2=\sec^2-1\]

OpenStudy (anonymous):

use 1+tan^2A=sec^2A AND cot^2+1=cosec^2

OpenStudy (anonymous):

So if we factor \[\Large (\sec^4-1)=(\sec^2-1)(\sec^2+1)\]

OpenStudy (anonymous):

Remembering that \[\Large \sec^2-1 =\tan^2\] We have \[\Large \tan^2(\sec^2+1)\]

OpenStudy (anonymous):

So now replace sec^2 with the identified before and tell me what you get

OpenStudy (anonymous):

the identity

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

(1+tan^2)+(1+tan^2)^4 this okay?

OpenStudy (anonymous):

Your notation is incorrect though

OpenStudy (anonymous):

\[\Large \tan^2((1+\tan^2)+1)\]

OpenStudy (anonymous):

So with that simplified, you can get your other side of the equation

OpenStudy (anonymous):

l.h.s sec^4A-1 =(sec^2A-1)(sec^2A+1) =tan^2A{(tan^2A+1)+1} am i okay rng agn?

OpenStudy (anonymous):

That's fine, the two 1s add together and then that's where you get the 2tan^2 and the tan^2s multply together and you get tan^4

OpenStudy (anonymous):

oh now =tan^2A(tan^2A+2) =tan^4a+2tan^2A =r.hs i think now okay

OpenStudy (anonymous):

thanks you doulikepiecauseidont.

OpenStudy (anonymous):

Yw. Anytime